### v1

# Learning sparsity-promoting regularizers for linear inverse problems

Abstract
This paper introduces a novel approach to learning sparsity-promoting regularizers for solving linear inverse problems. We develop a bilevel optimization framework to select an optimal synthesis operator, denoted as $B$, which regularizes the inverse problem while promoting sparsity in the solution. The method leverages statistical properties of the underlying data and incorporates prior knowledge through the choice of $B$. We establish the well-posedness of the optimization problem, provide theoretical guarantees for the learning process, and present sample complexity bounds. The approach is demonstrated through examples, including compact perturbations of a known operator and the problem of learning the mother wavelet, showcasing its flexibility in incorporating prior knowledge into the regularization framework. This work extends previous efforts in Tikhonov regularization by addressing non-differentiable norms and proposing a data-driven approach for sparse regularization in infinite dimensions.
Keywords: inverse problems, statistical learning, bilevel optimization, operator learning, sparsity-promoting regularization.
1 Problem formulation and main contributions
Consider a linear inverse problem
$y=Ax+\varepsilon,$ (1)
where we assume that $A\colon X\rightarrow Y$ is a linear bounded operator between the Hilbert spaces $X,Y$, whereas the inverse $A^{-1}$ (if it exists) can be in general an unbounded operator.
We introduce a variational strategy [20] to regularize the inverse problem (1), and together promote the sparsity of the solution with respect to a suitable basis or frame. To do so, we consider an operator $B\in\mathcal{L}(\ell^{2},X)$ (where $\mathcal{L}(\ell^{2},X)$ denotes the space of bounded linear operators from the sequence space $\ell^{2}(\mathbb{N})$ to the Hilbert space $X$, equipped with the operator norm) and define the following minimization problem:
$\hat{x}_{B}=B\hat{u}_{B},\quad\hat{u}_{B}=\operatorname*{arg\,min}_{u\in\ell^{2}}\Big{\{}\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}ABu\|_{Y}^{2}-\langle y,\Sigma_{\varepsilon}^{-1}ABu\rangle_{Y}+\|u\|_{\ell^{1}}\Big{\}},$ (2)
where $\Sigma_{\varepsilon}$ is the covariance of the noise $\varepsilon$ (see Assumption 1.1 below). In Section 2, we introduce a set of assumptions that guarantee the well-definedness and well-posedness of problem (2).
To better interpret the proposed regularization strategy, we remark that in a finite-dimensional setup (let, e.g., $X=\mathbb{R}^{n}$ and $Y\in\mathbb{R}^{m}$), if the matrix $B\in\mathbb{R}^{n\times n}$ is invertible, the minimization problem (2) admits the following formulation:
$\hat{x}_{B}=\operatorname*{arg\,min}_{x\in\mathbb{R}^{n}}\Big{\{}\frac{1}{2}\|Ax-y\|^{2}_{\Sigma_{\varepsilon}}+\|B^{-1}x\|_{1}\Big{\}},$
where $\|\cdot\|_{\Sigma_{\varepsilon}}=\|\Sigma_{\varepsilon}^{-1/2}\cdot\|$ is a norm on $\mathbb{R}^{m}$ that leverages the knowledge of the noise covariance $\Sigma_{\varepsilon}$ to whiten the residual error. Further details about this simplified formulation can be found in Appendix A.
The main focus of this paper is related to the choice of the synthesis operator $B\colon\ell^{2}\rightarrow X$ within (2). In particular, let us introduce the map
[…]
For a fixed $B$ satisfying the theoretical requirements expressed in Section 2, the map $R_{B}$ is a stable reconstruction operator, i.e., it provides a continuous approximation of the inverse map $A^{-1}$. In particular, $R_{B}$ promotes the sparsity of the regularized solution in terms of the synthesis operator $B$: namely, $R_{B}(y)=B\hat{u}_{B}$ is chosen so that it balances a good data fidelity ($AR_{B}(y)\approx y$) and only a few components of $\hat{u}_{B}$ are different from 00.
The choice of the synthesis operator $B$ encodes crucial information regarding the prior distribution of $x$. Consider, for simplicity, a signal processing problem (such as denoising, or deblurring), which, after a discretization of the spaces of signals, can be formulated as a linear inverse problem in $X=Y=\mathbb{R}^{n}$. Then, promoting the sparsity of the reconstructed signals under the choice $B=\operatorname{Id}$ encodes the prior information that the ground truths are expected to have few, isolated, spikes. Choosing $B$ to be a basis of (discretized) sines and cosines promotes band-limited signals in the Fourier domain. Setting $B$ to a wavelet transform, instead, might promote signals showing few jump discontinuities and smooth anywhere else. In general, the synthesis operator $B$ should be carefully chosen to incorporate, in the reconstruction operator $R_{B}$, any significant prior knowledge, or prior belief, on the ground truths.
In this paper, we follow a statistical learning approach for the selection of $B$. In particular, we assume that $x$ and $y$ are random objects with a joint probability distribution $\rho$. We then leverage the (partial) knowledge of such a probability distribution to define a data-driven rule for selecting $B$. To do so, we must introduce some assumptions on the probability distribution $\rho$, which are carefully detailed in Section 3. Before delving into the details, though, let us briefly sketch the overall idea of the learning-based choice of $B$, introducing some minimal requirements on the random objects $x,y$.
Assumption 1.1.
Let $(x,y)\sim\rho$, being $y=Ax+\varepsilon$, and where:
  • $x$ is a square-integrable random variable on $X$;
  • $\varepsilon$ is a zero-mean square-integrable random variable on $Y$, independent of $x$, with known, trace-class, and injective covariance $\Sigma_{\varepsilon}\colon Y\rightarrow Y$.
The assumption that $\varepsilon$ is a zero-mean random variable with $\mathbb{E}{[\left\|\varepsilon\right\|^{2}_{Y}]}<+\infty$ is in principle quite restrictive and, for example, the case of white noise requires careful treatment when $Y$ is infinite-dimensional. In its most common formulation, white noise is modeled as a random process on a Hilbert space (for example, $Y=L^{2}(\Omega)$) with zero mean and covariance equal to the identity. Such a noise model clearly fails in satisfying the hypotheses whenever $Y$ is infinite dimensional, because the covariance $\Sigma_{\varepsilon}=\operatorname{Id}$ is not trace-class, but only bounded. However, it is possible to represent such a process as a square-integrable random variable by carefully selecting the space $Y$, as described in Appendix B.
For a specific choice of synthesis operator $B$, the quality of the reconstruction operator $R_{B}$ can be evaluated through the expected loss $L\colon\mathcal{L}(\ell^{2},X)\rightarrow\mathbb{R}$:
$L(B)=\mathbb{E}_{(x,y)\sim\rho}[\|R_{B}(y)-x\|_{X}^{2}].$ (4)
Let us now consider a suitable class of operators $\mathcal{B}\subset\mathcal{L}(\ell^{2},X)$: we define the optimal regularizer $R^{\star}=R_{B^{\star}}$, where $B^{\star}$ is a minimizer of the expected loss over the set of admissible operators $\mathcal{B}$. In particular, putting together the expressions of (4) and (2), we obtain:
$\begin{gathered}B^{\star}\in\operatorname*{arg\,min}_{B\in\mathcal{B}}L(B),\\ R_{B}(y)=B\operatorname*{arg\,min}_{u\in\ell^{2}}\Big{\{}\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}ABu\|_{Y}^{2}-\langle y,\Sigma_{\varepsilon}^{-1}ABu\rangle_{Y}+\|u\|_{\ell^{1}}\Big{\}},\end{gathered}$ (5)
which is usually referred to as a bilevel optimization problem. Note that the optimality of $B^{\star}$ strongly depends on the choice of the class $\mathcal{B}$.
We moreover stress that the optimal target $B^{\star}$ can only be computed if the joint probability distribution $\rho$ of $x$ and $y$ is known. This is of course not verified in practical applications, but in many contexts, we may suppose to have access to a sample of $m$ pairs ${\mathbf{z}}=\{(x_{j},y_{j})\}_{j=1}^{m}$ such that the family is independent and identically distributed as $(x,y)$. In this case, following the paradigm of supervised learning, we can approximate the expected loss by an empirical average, also known as empirical risk, namely
$\widehat{L}(B)=\frac{1}{m}\sum_{j=1}^{m}\|R_{B}(y_{j})-x_{j}\|_{X}^{2}.$ (6)
A natural estimator of $B^{\star}$ is then given by any minimizer $\widehat{B}$ of the empirical loss:
$\widehat{B}\in\operatorname*{arg\,min}_{B\in\mathcal{B}}\widehat{L}(B).$ (7)
We stress that both $\widehat{L}(B)$ and $\widehat{B}$ are random variables depending on the sample ${\mathbf{z}}$. We simply denote this dependence by $\widehat{\cdot}$. The task of sample error estimates is to quantify the dependence of the empirical target $\widehat{B}$ on the sample ${\mathbf{z}}$, in particular by bounding (either in probability or in expectation) the excess risk $L(\widehat{B})-L(B^{\star})$ in terms of the sample size $m$.
Main contribution of this paper
In our analysis, we pursue the following main goals:
  1. 1.
    studying the theoretical properties of the inner minimization problem (2), and in particular its well-posedness for fixed $B$ and the continuous dependence of the minimizer $\hat{x}_{B}$ in terms of $B\in\mathcal{B}$ (see Theorems 2.3 and 2.5 in Section 2);
  2. 2.
    studying the approximation properties of the empirical target $\widehat{B}$ (constructed leveraging the knowledge of $A$, $\Sigma_{\varepsilon}$, and of the sample ${\mathbf{z}}$) with respect to the optimal target $B^{\star}$, deriving sample error estimates for the excess risk (see Theorem 3.2 in Section 3);
  3. 3.
    formulate some relevant applications which satisfy the assumptions on $x$, $y$, $A$, and $\mathcal{B}$ introduced in Sections 2 and 3 (see Section 4).
Comparison with existing literature
This work originated from the analysis carried out by a subset of the authors in [4], where we considered the problem of learning (also) the optimal operator $B$ in generalized Tikhonov regularization, i.e., when a quadratic penalty term is considered within the inner problem (2) instead of the $\ell^{1}$ norm. The main difficulty associated with the proposed extensions resides in the lack of strong convexity and differentiability of the $\ell^{1}$ norm. Moreover, unlike in the Tikhonov case, the inner minimization problem (2) does not possess a closed-form solution: unfortunately, this also prevents the explicit computation of the optimal regularizer $B^{\star}$, which was one of the main results in [4]. As a final consequence, it is not possible to formulate here an unsupervised strategy to learn $B^{\star}$, i.e., based only on a training set of ground truths $\{x_{j}\}_{j=1}^{m}$: indeed, the unsupervised strategy proposed in [4] extensively leveraged the explicit expression of the optimal operator $B^{\star}$.
Other extensions of the work [4] have been carried out in [38, 12, 15, 9, 5] even though a statistical learning approach for sparse optimization in infinite dimension was not considered yet. See also [24] for a connection between generalized Tikhonov and linear plug-and-play denoisers. Let us just mention that bilevel approaches for inverse problems in imaging have been studied, from numerical and optimization points of view since many years [13, 19]. In particular, the works [26, 27, 22, 21] study the bilevel learning of $\ell^{1}$ regularizers, focusing on the finite-dimensional case and mostly on the algorithmic aspects, without considering the generalization issue from the theoretical point of view. Moreover, our work is also related to statistical inverse learning for inverse problems [7, 25, 11] and, also, more generally to the growing field of operator learning [36, 31, 30, 18, 8]. Let us also mention some of the main works on using machine learning techniques for solving inverse problems that had motivated this work [2, 29, 3, 33, 6, 32, 34, 35].
It is worth observing that the problem of learning an operator $B$ yielding a sparse representation of a dataset $\{x_{j}\}_{j=1}^{m}$ is deeply connected with the well-known task of Dictionary Learning. Although it is possible to provide a formulation of a dictionary learning problem very close to the bilevel problem (5), the two problems pursue two distinct aims and may lead to different results, as shown in Appendix C. Despite this, the sample complexity bounds we obtain in this work are comparable to those derived for dictionary learning [23, 39].
2 Theoretical results on the deterministic optimization problem (2)
The goal of this section is to study the well-posedness of the minimization problem formulated in (2) for a fixed $B\in\mathcal{L}(\ell^{2},X)$. Moreover, we study the stability of the minimization with respect to perturbations of $B$. To do so, let us introduce the following set of hypotheses.
Assumption 2.1 (Compatibility assumption between $\Sigma_{\varepsilon}$ and $A$).
Let the covariance matrix $\Sigma_{\varepsilon}$ satisfy Assumption 1.1, and assume $\operatorname{Im}(A)\subset\operatorname{Im}(\Sigma_{\varepsilon})$. Moreover, we assume that
$\Sigma_{\varepsilon}^{-1}A\colon X\rightarrow Y\text{ is a compact operator}.$ (8)
For the rest of the paper, we use the convention $J_{B}(u)=F_{B}(u)+\Phi(u)$, where
$F_{B}(u)=\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}ABu\|_{Y}^{2}-\langle y,\Sigma_{\varepsilon}^{-1}ABu\rangle_{Y}\quad\text{and}\quad\Phi(u)=\|u\|_{\ell^{1}}.$ (9)
Notice that, thanks to Assumption 2.1, the functional $J_{B}$ is well-defined for fixed $B\in\mathcal{L}(\ell^{2},X)$. Indeed, the first term can be rewritten as $\frac{1}{2}\langle\Sigma_{\varepsilon}^{-1}ABu,ABu\rangle_{Y}$, and $\Sigma_{\varepsilon}^{-1}ABu$ exists and is unique since $\operatorname{Im}(AB)\subset\operatorname{Im}\Sigma_{\varepsilon}=\operatorname{dom}(\Sigma_{\varepsilon}^{-1})$, and analogously for the second term. Finally, since the minimization problem (2) is set in the Hilbert space $\ell^{2}\supset\ell^{1}$, the third term in (2) should be interpreted as $\|u\|_{\ell^{1}}$ if $u\in\ell^{1}$ and +$\infty$+ ∞ if $u\in\ell^{2}\setminus\ell^{1}$.
Since $J_{B}$ is non-differentiable and not strictly convex, further assumptions on the interplay between $A$ and $B$ are needed to guarantee well-posedness of the minimization task. Our analysis is confined to the case where $AB$ is assumed to be finite basis injective.
Assumption 2.2 (Finite Basis Injectivity (FBI) of $AB$).
For all $I\subset\mathbb{N}$ with $\operatorname{card}(I)<\infty$, the operator $(AB)|_{\ell^{2}_{I}}$ is injective, where we denoted =$\mathbb{N}=\{1,2,\dots\}$ and $\ell^{2}_{I}:=\operatorname{span}\{e_{i}:i\in I\}$.
Note that while this assumption is satisfied when $B$ is injective(provided $A$ is injective), it also holds for more general structures (e.g. FBI frames). The well-posedness of the minimization problem (2) is now characterized by the following theorem.
Theorem 2.3.
Let $A\in\mathcal{L}(X,Y),\Sigma_{\varepsilon}\in\mathcal{L}(Y,Y)$, and $B\in\mathcal{L}(\ell^{2},X)$ satisfy Assumptions 2.1 and 2.2, and let $y\in Y$. There exists a unique minimizer $\hat{u}_{B}=\hat{u}_{B}(y)$ of $J_{B}$, and, consequently, a unique $\hat{x}_{B}=B\hat{u}_{B}$ in (2).
The proof is presented in two parts: the existence is proved at the end of Section 2.1 while the uniqueness follows as a consequence of Theorem 2.5 at the end of Section 2.2.
Let us next consider the stability of the minimizer with respect to perturbing $B$ in a suitable class of operators $\mathcal{B}$. Towards that end, let us introduce the following assumption.
Assumption 2.4 (Requirements on $\mathcal{B}$).
Let $\mathcal{B}\subset\mathcal{L}(\ell^{2},X)$ be a compact set of operators such that every $B\in\mathcal{B}$ satisfies Assumption 2.2.
A first consequence of the compactness of $\mathcal{B}$ is that the norms $\|B\|$, $B\in\mathcal{B}$, are uniformly bounded by $|{\mathcal{B}}|:=\sup_{B\in{\mathcal{B}}}\left\|B\right\|<\infty$. Moreover, as we show in the following subsections, Assumption 2.4 allows us to provide a uniform expression of several properties of the solutions of problem (2), namely, independently of the choice of $B$. The most relevant consequence of this discussion is the following result, providing a global stability estimate of the minimizers $\hat{x}_{B}$ with respect to perturbations of $B$ within $\mathcal{B}$.
Theorem 2.5.
Let $\|y\|_{Y}\leq Q.$ There exists a constant $c_{\rm ST}=c_{\rm ST}(\mathcal{B},A,\Sigma_{\varepsilon},Q)$ such that for every $B_{1},B_{2}\in\mathcal{B}$,
$\|R_{B_{1}}(y)-R_{B_{2}}(y)\|_{X}\leq c_{\rm ST}\|B_{1}-B_{2}\|^{1/2},$ (10)
2.1 Existence of the minimizer with fixed $B\in{\mathcal{B}}$
To establish the existence of a minimizer $\hat{u}_{B}$ in (2), we first show that the sublevel sets of $J_{B}$ are bounded, which ensures the weak convergence of any minimizing sequence.
Lemma 2.6.
There exists a monotonically increasing function $h\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ depending on $A$, $\Sigma_{\varepsilon}$, and $|\mathcal{B}|$ such that the following holds: if $\left\|w\right\|_{X}\leq|{\mathcal{B}}|$ and $\left\|\Sigma_{\varepsilon}^{-1}Aw\right\|_{Y}\geq\gamma>0$, then $\left\|\Sigma_{\varepsilon}^{-1/2}Aw\right\|_{Y}\geq h(\gamma)>0$.
Proof.
Let us prove existence of a general $h$ by contradiction: assume that for all $n$ there exists $w_{n}\in X$ satisfying $\left\|w_{n}\right\|_{X}\leq|{\mathcal{B}}|$, $\left\|\Sigma_{\varepsilon}^{-1}Aw_{n}\right\|_{Y}\geq\gamma$ and $\left\|\Sigma_{\varepsilon}^{-1/2}Aw_{n}\right\|_{Y}<\frac{1}{n}$. Since $\{w_{n}\}_{n=1}^{\infty}\subset X$ is bounded, we can find a subsequence $\{w_{n_{k}}\}_{k=1}^{\infty}$ that weakly converges to some $w\in X$. Due to weak lower semicontinuity of the norm, we must have $w\in\operatorname{ker}(A)$ and, therefore, by the compactness of $\Sigma_{\varepsilon}^{-1}A$ we also have that $\|\Sigma_{\varepsilon}^{-1}Aw_{n}\|_{Y}\rightarrow 0$. This contradicts with our assumption and, therefore, a lower bound $h(\gamma)>0$ must exist.
The existence of a monotonically increasing $h$ can also be demonstrated by contradiction: suppose that there exist $\gamma_{1}<\gamma_{2}$ such that for any $h$ satisfying the statement, we have $h(\gamma_{1})>h(\gamma_{2})$. Then we immediately see that $\tilde{h}$ defined by
$\tilde{h}(\gamma)=\begin{cases}h(\gamma)&\gamma\neq\gamma_{1}\\ h(\gamma_{2})&\gamma=\gamma_{1},\end{cases}$
also satisfies the claim yielding the contradiction. ∎
Proposition 2.7.
Let $M\in\mathbb{R}$, and suppose Assumptions 1.1, 2.1, 2.2 and 2.4 hold. Let $\|y\|_{Y}\leq Q$. Then, there exists a constant $c_{\rm UB}(M)=c_{\rm UB}(M;A,\Sigma_{\varepsilon},Q,|{\mathcal{B}}|)$ such that, for every $B\in\mathcal{B}$, we have
$\{u\in\ell^{2}\;|\;J_{B}(u)\leq M\}\subset\{u\in\ell^{2}\;|\;\|u\|_{\ell^{1}}\leq c_{\rm UB}(M)\}.$ (11)
Proof.
The condition $J_{B}(u)\leq M$ reads as
$\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}ABu\|_{Y}^{2}+\left\|u\right\|_{\ell^{1}}\leq M+\langle y,\Sigma_{\varepsilon}^{-1}ABu\rangle_{Y}.$
Let us next make a change of variables with $w=Bu/\left\|u\right\|_{\ell^{1}}$ and $\tau=\left\|u\right\|_{\ell^{1}}$, which leads to
$\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}Aw\|_{Y}^{2}\tau^{2}+\tau\leq M+\tau\left\|y\right\|_{Y}\left\|\Sigma_{\varepsilon}^{-1}Aw\right\|_{Y}.$ (12)
In particular, we have
$1\leq\frac{M}{\tau}+\left\|y\right\|_{Y}\left\|\Sigma_{\varepsilon}^{-1}Aw\right\|_{Y}.$
Hence, either $\left\|u\right\|_{\ell^{1}}=\tau\leq\max\{2M,0\}$ or for any $\tau\geq\max\{2M,0\}$ it holds that
$\left\|\Sigma_{\varepsilon}^{-1}Aw\right\|_{Y}\geq\frac{1}{2\left\|y\right\|_{Y}}.$
Since $\left\|Bu\right\|_{X}\leq|\mathcal{B}|\left\|u\right\|_{\ell^{2}}$ for all $u\in\ell^{2}$ and $B\in\mathcal{B}$, we have
$\|u\|_{\ell^{1}}\geq\|u\|_{\ell^{2}}\geq\frac{1}{|{\mathcal{B}|}}\|Bu\|_{X}\qquad\forall u\in\ell^{2},\forall B\in\mathcal{B},$
which also implies that $\left\|w\right\|_{X}\leq|{\mathcal{B}}|$. Next, by Lemma 2.6 it follows that
$\left\|\Sigma_{\varepsilon}^{-1/2}Aw\right\|_{Y}\geq h,$
where we abbreviate $h=h\left(\frac{1}{2\left\|y\right\|_{Y}}\right)$ for convenience. Modifying (12) we obtain
$\frac{1}{2}h^{2}\tau^{2}\leq M+\tau\left\|y\right\|_{Y}\left\|\Sigma_{\varepsilon}^{-1}A\right\||\mathcal{B}|\leq M+\frac{1}{4}h^{2}\tau^{2}+\frac{\left\|y\right\|_{Y}^{2}\left\|\Sigma_{\varepsilon}^{-1}A\right\|^{2}|\mathcal{B}|^{2}}{h^{2}}$
and solving for $\tau$ yields
$\tau^{2}\leq\frac{4M}{h^{2}}+\frac{4\left\|y\right\|_{Y}^{2}\left\|\Sigma_{\varepsilon}^{-1}A\right\|^{2}|\mathcal{B}|^{2}}{h^{4}}.$ (13)
Note carefully that, by the monotonicity of $h$, from $\|y\|_{Y}\leq Q$ it follows that $h\geq h\left(\frac{1}{2Q}\right)$, hence
$\tau^{2}\leq\frac{4M}{h\left(\frac{1}{2Q}\right)^{2}}+\frac{4Q^{2}\left\|\Sigma_{\varepsilon}^{-1}A\right\|^{2}|\mathcal{B}|^{2}}{h\left(\frac{1}{2Q}\right)^{4}}.$ (14)
Now the desired bound for $\tau=\left\|u\right\|_{\ell^{1}}$ follows for any $u\in\ell^{2}$, $B\in\mathcal{B}$. ∎
Proof of Thm. 2.3 (existence).
The existence of a minimizer $\hat{u}_{B}$ in (2) is now guaranteed by standard arguments: any minimizing sequence $\{u_{j}\}_{j=1}^{\infty}\subset\ell^{2}$ belongs to some sublevel set and, therefore, to some $\ell^{1}$-ball according to Proposition 2.7. By Banach–Alaoglu theorem, the sequence has a weak-* converging subsequence in $\ell^{1}$. Since $F_{B}$ is continuous and the $\ell^{1}$-norm is lower semicontinuous with respect to the weak-* topology, one can show that the limit is a minimizer of $J_{B}$. ∎
Let us note that since $J_{B}(\hat{u}_{B})\leq J_{B}(0)=0$, denoting by $c_{\rm UB}$ the constant $c_{\rm UB}(0)$ in (11) we have the following bound for such minimizers, uniformly in $B\in\mathcal{B}$.
$\|\hat{u}_{B}\|_{\ell^{2}}\leq\|\hat{u}_{B}\|_{\ell^{1}}\leq c_{\rm UB}.$ (15)
The uniqueness of such a minimizer is a consequence of Proposition 2.10, as we will show later.
2.2 Stability in ${\mathcal{B}}$ and uniqueness of the minimizer
The key result of this section is Proposition 2.10, as it directly enables the uniqueness of the minimizer in Theorem 2.3, and its stability with respect to $B$. Such a result is in the spirit of [10, Theorem 2], and its proof is partially based on the one of [10, Lemma 3]. The main contribution of Proposition 2.10 is that the constant appearing in (19) does not depend on the choice of $B$, provided that we consider an operator class ${\mathcal{B}}$ satisfying Assumption 2.4. The proof of Proposition 2.10 requires some preliminary lemmas.
Let us first note that $F_{B}\colon\ell^{2}\rightarrow\mathbb{R}$ is convex and differentiable with gradient
$F_{B}^{\prime}(u)=B^{*}\left(\Sigma_{\varepsilon}^{-1}A\right)^{*}(ABu-y),$
where we have used the identity $A^{*}(\Sigma_{\varepsilon}^{-1}A)=\left(\Sigma_{\varepsilon}^{-1}A\right)^{*}A$, since $\langle\Sigma_{\varepsilon}^{-1}Av,Aw\rangle=\langle Av,\Sigma_{\varepsilon}^{-1}Aw\rangle=\langle(\Sigma_{\varepsilon}^{-1}A)^{*}Av,w\rangle$ for all $v,w\in X$ as $\Sigma_{\varepsilon}^{-1}$ is self-adjoint and $\operatorname{Im}(A)\subset\operatorname{dom}({\Sigma_{\varepsilon}^{-1}})$. Notice that $F_{B}^{\prime}$ is affine in $u$ and Lipschitz continuous, and the Lipschitz constant is uniformly bounded by $L_{F^{\prime}}=|{\mathcal{B}}|^{2}\|\Sigma_{\varepsilon}^{-1/2}A\|^{2}$ for $B\in{\mathcal{B}}$. Since both $F_{B}$ and $\Phi$ are convex, $\hat{u}_{B}$ being a minimizer of $J_{B}$ is equivalent to 00 belonging to the subgradient set of $J_{B}$ at $\hat{u}_{B}$, i.e. $0\in\partial J_{B}(\hat{u}_{B})$. Now, due to the differentiability of $F_{B}$, an equivalent optimality condition is given by $\hat{w}_{B}\in\partial\Phi(\hat{u}_{B})$, where
$\hat{w}_{B}=-F^{\prime}_{B}(\hat{u}_{B})=B^{*}\left(\Sigma_{\varepsilon}^{-1}A\right)^{*}(y-AB\hat{u}_{B}).$ (16)
Moreover, thanks to the explicit expression of the subdifferential of the $\ell^{1}$-norm, we can specify that $\hat{w}_{B}\in\partial\Phi(\hat{u}_{B})$ is equivalent to
$\hat{w}_{B,k}=\left\{\begin{aligned} \mathrm{sign}(\hat{u}_{B,k})\quad&\text{if }\hat{u}_{B,k}\neq 0\\ \xi\in[-1,1]\quad&\text{if }\hat{u}_{B,k}=0.\\ \end{aligned}\right.$ (17)
The following lemmas explore properties that are valid uniformly in the compact operator class $\mathcal{B}$. For $N>0$, we introduce the orthogonal projection operator onto $\ell^{2}$ by setting
$P_{N}\colon\ell^{2}\rightarrow\ell^{2}:\ P_{N}u=(u_{1},\ldots,u_{N},0,0,\ldots),\qquad P_{N}^{\perp}=\operatorname{Id}-P_{N}.$
Lemma 2.8.
Suppose that Assumptions 2.1, 2.2, and 2.4 are satisfied. Let $\mathbb{B}_{Y}(Q)=\{y\in Y:\|y\|_{Y}\leq Q\}$. Then the elements $\hat{w}_{B}\in\ell^{2}$ corresponding to the minimizers $\hat{u}_{B}\in\ell^{2}$ of $J_{B}$ via (16) satisfy
$\lim_{N\rightarrow\infty}\sup\big{\{}\|P_{N}^{\perp}\hat{w}_{B}\|_{\ell^{2}}:\;B\in\mathcal{B},\ y\in\mathbb{B}_{Y}(Q)\big{\}}=0.$
Proof.
By contradiction, suppose there exist $\varepsilon>0$ and a diverging sequence $N_{M}$, $M=1,2,...$, such that
$\sup\bigl{\{}\big{\|}P_{N_{M}}^{\perp}B^{*}\big{(}\Sigma_{\varepsilon}^{-1}A\big{)}^{*}\big{(}y-AB\hat{u}_{B}\big{)}\big{\|}_{\ell^{2}}:B\in\mathcal{B},\ y\in\mathbb{B}_{Y}(Q)\bigr{\}}\geq 2\varepsilon\quad\forall M.$
Consider in particular a sequence $\{B_{M}\}_{M=1}^{\infty}\subset\mathcal{B}$ and a sequence $\{y_{M}\}_{M=1}^{\infty}\subset\mathbb{B}_{Y}(Q)$ satisfying
$\big{\|}P_{N_{M}}^{\perp}B_{M}^{*}\big{(}\Sigma_{\varepsilon}^{-1}A\big{)}^{*}\big{(}y_{M}-AB_{M}\hat{u}_{B_{M}}\big{)}\big{\|}_{\ell^{2}}\geq\varepsilon\quad\forall M,$
being $\hat{u}_{B_{M}}$ the solution of (2) with $B=B_{M}$ and $y=y_{M}$. By the compactness of $\mathcal{B}$, there exists $B\in\mathcal{B}$ such that $B_{M}\rightarrow B$ in the strong operator topology, up to a subsequence. Moreover, consider the sequence $h_{M}=AB_{M}\hat{u}_{B_{M}}$, $M=1,2,...$. Since $\left\|\hat{u}_{B_{M}}\right\|$ is bounded uniformly with respect to $B\in\mathcal{B}$ and $y\in\mathbb{B}_{Y}(Q)$ by (15), $\left\|B_{M}\right\|\leq|{\mathcal{B}}|$ independently of $M$, and $A$ is bounded, the sequence $\{h_{M}\}_{M=1}^{\infty}\subset Y$ is bounded, thus weakly convergent (up to a subsequence) to an element $h\in Y$. As a consequence of the compactness of $\big{(}\Sigma_{\varepsilon}^{-1}A\big{)}^{*}$ by the Schauder theorem, we have that $(\Sigma_{\varepsilon}^{-1}A)^{*}h_{M}\rightarrow(\Sigma_{\varepsilon}^{-1}A)^{*}h$ in $X$. Similarly, the bounded sequence $\{y_{M}\}_{M=1}^{\infty}$ admits a weak limit $y\in Y$, and $(\Sigma_{\varepsilon}^{-1}A)^{*}y_{M}\rightarrow(\Sigma_{\varepsilon}^{-1}A)^{*}y$.
In conclusion, we have that
$\varepsilon$$\leq$ $\|P_{N_{M}}^{\perp}B_{M}^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(y_{M}-h_{M})\|_{\ell^{2}}$
$\leq$ $\|P_{N_{M}}^{\perp}B_{M}^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(y_{M}-y)\|_{\ell^{2}}+\|P_{N_{M}}^{\perp}B_{M}^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(h_{M}-h)\|_{\ell^{2}}$
$\quad+\|P_{N_{M}}^{\perp}B_{M}^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(y-h)\|_{\ell^{2}}$
$\leq$ $|\mathcal{B}|\|(\Sigma_{\varepsilon}^{-1}A)^{*}(y_{M}-y)\|_{X}+|\mathcal{B}|\|(\Sigma_{\varepsilon}^{-1}A)^{*}(h_{M}-h)\|_{X}$
$\quad+\|P_{N_{M}}^{\perp}(B_{M}-B)^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(y-h)\|_{\ell^{2}}+\|P_{N_{M}}^{\perp}B^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(y-h)\|_{\ell^{2}}$
$\leq$ $|\mathcal{B}|\|(\Sigma_{\varepsilon}^{-1}A)^{*}(y_{M}-y)\|_{X}+|\mathcal{B}|\|(\Sigma_{\varepsilon}^{-1}A)^{*}(h_{M}-h)\|_{X}$
$\quad+\|B_{M}-B\|_{\ell^{2}\rightarrow X}\|(\Sigma_{\varepsilon}^{-1}A)^{*}(y-h)\|_{X}+\|P_{N_{M}}^{\perp}B^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(y-h)\|_{\ell^{2}}.$
As $M\rightarrow\infty$, all the terms on the right-hand side converge to 00, which entails a contradiction. ∎
Lemma 2.9.
Suppose that Assumptions 2.1, 2.2, and 2.4 are satisfied. Take $N\in\mathbb{N}$. Then, there exists a constant $c_{\rm LB}=c_{\rm LB}(A,\mathcal{B},N)>0$ such that
$\|\Sigma_{\varepsilon}^{-1/2}ABP_{N}u\|_{Y}^{2}\geq c_{\rm LB}\|P_{N}u\|_{\ell^{2}}^{2}$
for all $B\in{\mathcal{B}}$ and $u\in\ell^{2}$.
Proof.
We prove this claim by contradiction. Assume that there exist two sequences $\{u_{M}\}_{M=1}^{\infty}\subset\ell^{2}$ and $\{B_{M}\}_{M=1}^{\infty}\subset\mathcal{B}$ such that
$\frac{\|\Sigma_{\varepsilon}^{-1/2}AB_{M}P_{N}u_{M}\|_{Y}^{2}}{\|P_{N}u_{M}\|_{\ell^{2}}^{2}}<\frac{1}{M}\quad\forall M.$ (18)
Let us write $p_{M}=\frac{P_{N}u_{M}}{\|P_{N}u_{M}\|_{\ell^{2}}}$ to obtain that a sequence $\{p_{M}\}_{M=1}^{\infty}\subset H_{N}$, where $H_{N}=\{u\in\ell^{2}:u_{k}=0\ \forall k>N,\|u\|_{\ell^{2}}=1\}$. Notice that $H_{N}\subset X$ is a finite-dimensional, bounded and closed subset and hence compact. Rephrasing (18) we have
$\|\Sigma_{\varepsilon}^{-1/2}AB_{M}p_{M}\|_{Y}^{2}<\frac{1}{M}\quad\forall M.$
Since both $\{p_{M}\}_{M=1}^{\infty}\subset H_{N}$ and $\{B_{M}\}_{M=1}^{\infty}\subset{\mathcal{B}}$ belong to compact sets, up to a subsequence, the sequences converge to limit points $p\in H$ and $B\in\mathcal{B}$, respectively. By the continuity of $\Sigma_{\varepsilon}^{-1/2}A$ and by the equi-continuity of the operators in $\mathcal{B}$, we deduce that $\|\Sigma_{\varepsilon}^{-1/2}ABp\|_{Y}=0$, which by the injectivity of $\Sigma_{\varepsilon}^{-1/2}$ and by Assumption 2.2 implies that $p=0$. This yields a contradiction with the assumption that $\|p\|=1$ and completes the proof. ∎
We can finally state and prove the main result of this subsection:
Proposition 2.10.
Let $\hat{u}_{B}\in\ell^{2}$ be a minimizer of $J_{B}$ being $B\in\mathcal{B}$ and $y\in\mathbb{B}_{Y}(Q)$. For every $M$, there exists a constant $\tilde{c}_{\rm ST}=\tilde{c}_{\rm ST}(M;A,\Sigma_{\varepsilon},\mathcal{B},Q)$ such that for any $B\in{\mathcal{B}}$ we have
$\{v\in\ell^{2}\;|\;J_{B}(v)\leq M\}\subset\{v\in\ell^{2}\;|\;\|v-\hat{u}_{B}\|_{\ell^{2}}^{2}\leq\tilde{c}_{\rm ST}(J_{B}(v)-J_{B}(\hat{u}_{B}))\}.$ (19)
Proof.
For clarity, the proof is divided into several steps.
Step $1$. Let us show that there exists $N_{0}=N_{0}(A,\Sigma_{\varepsilon},\mathcal{B},Q)$ such that $P_{N_{0}}^{\perp}\hat{u}_{B}=0$. Rephrasing Lemma 2.8, we have that there exists a sequence $\delta_{N}$, $N=1,2,...,$ converging to zero such that
$\|P_{N}^{\perp}\hat{w}_{B}\|_{\ell^{2}}\leq\delta_{N}\quad\forall B\in\mathcal{B},\ y\in\mathbb{B}_{Y}(Q).$
Therefore, it is possible to pick $N_{0}$ such that
$\|P_{N_{0}}^{\perp}\hat{w}_{B}\|_{\ell^{2}}\leq\frac{1}{2}$
for all $B\in{\mathcal{B}}$ and $y\in\mathbb{B}_{Y}(Q)$. Notice that such $N_{0}$ is independent of the specific choice of $B$, and depends on $y$ only through $Q$: thus, we denote it as $N_{0}(A,\Sigma_{\varepsilon},\mathcal{B},Q)$. Since $\|P_{N_{0}}^{\perp}\hat{w}_{B}\|_{\ell^{\infty}}\leq\|P_{N_{0}}^{\perp}\hat{w}_{B}\|_{\ell^{2}}$, we also deduce that
$|\hat{w}_{B,k}|:=\left|[\hat{w}_{B}]_{k}\right|\leq\frac{1}{2}\quad\text{for $k>{N_{0}}$,}$ (20)
and by the optimality condition satisfied by $\hat{w}_{B}$, together with the characterization of the subdifferential in (17) we get
$\hat{u}_{B,k}:=[\hat{u}_{B}]_{k}=0\quad\text{for $k>{N_{0}}$,}$ (21)
whence $P_{N_{0}}^{\perp}\hat{u}_{B}=0$, or $P_{N_{0}}\hat{u}_{B}=\hat{u}_{B}$, independently on $B$.
Step $2$. We show that
$\|P_{N_{0}}^{\perp}(v-\hat{u}_{B})\|_{\ell^{2}}^{2}\leq 2c_{\rm UB}(M)\left(J_{B}(v)-J_{B}(\hat{u}_{B})-\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\|_{Y}^{2}\right),$ (22)
where $c_{\rm UB}(M)=c_{\rm UB}(M;A,\Sigma_{\varepsilon},Q,|{\mathcal{B}}|)$ is given in Proposition 2.7. For the ease of notation, let us denote
$r_{B}(v)=J_{B}(v)-J_{B}(\hat{u}_{B})-\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\|_{Y}^{2}.$
By direct computations, we have
$r_{B}(v)$ $=\Phi(v)-\Phi(\hat{u}_{B})+\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}ABv\|_{Y}^{2}-\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}AB\hat{u}_{B}\|_{Y}^{2}-\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\|_{Y}^{2}$
$\qquad-\langle y,\Sigma_{\varepsilon}^{-1}AB(v-\hat{u}_{B})\rangle_{Y}$
$=\Phi(v)-\Phi(\hat{u}_{B})+\langle\Sigma_{\varepsilon}^{-1/2}AB\hat{u}_{B},\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\rangle_{Y}-\langle B^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}y,v-\hat{u}_{B}\rangle$
$=\Phi(v)-\Phi(\hat{u}_{B})+\langle B^{*}(\Sigma_{\varepsilon}^{-1}A)^{*}(AB\hat{u}_{B}-y),v-\hat{u}_{B}\rangle.$
Using the definition of $\hat{w}_{B}$ in (16), we obtain the expression
$r_{B}(v)=\Phi(v)-\Phi(\hat{u}_{B})-\langle\hat{w}_{B},v-\hat{u}_{B}\rangle=\sum_{k\in\mathbb{N}}\left(|v_{k}|-|\hat{u}_{B,k}|-\hat{w}_{B,k}(v_{k}-\hat{u}_{B,k})\right).$
Moreover, thanks to (17), it is easy to verify that each term of the previous summation is non-negative. Consider now $N_{0}$ introduced in Step 1: by (20) and (21),
$r_{B}(v)$ $\geq\sum_{k>{N_{0}}}\left(|v_{k}|-|\hat{u}_{B,k}|-\hat{w}_{B,k}(v_{k}-\hat{u}_{B,k})\right)=\sum_{k>{N_{0}}}\left(|v_{k}|-\hat{w}_{B,k}v_{k}\right)$
$\geq\sum_{k>{N_{0}}}\left(|v_{k}|-\frac{1}{2}|v_{k}|\right)=\frac{1}{2}\|P_{N_{0}}^{\perp}v\|_{\ell^{1}}\geq\frac{1}{2}\|P_{N_{0}}^{\perp}v\|_{\ell^{2}}=\frac{1}{2}\|P_{N_{0}}^{\perp}(v-\hat{u}_{B})\|_{\ell^{2}}.$
In conclusion,
$\|P_{N_{0}}^{\perp}(v-\hat{u}_{B})\|_{\ell^{2}}^{2}\leq 2r_{B}(v)\|P_{N_{0}}^{\perp}(v-\hat{u}_{B})\|_{\ell^{2}}\leq 2r_{B}(v)\|v\|_{\ell^{2}}.$
Since $J_{B}(v)\leq M$, by Proposition 2.7 we have $\|v\|_{\ell^{2}}\leq c_{\rm UB}(M)$, hence (22) holds.
Step $3$. Consider again the projection operators associated with the choice $N=N_{0}$ discussed in Step 1:
$\|v-\hat{u}_{B}\|_{\ell^{2}}^{2}=\|P_{N_{0}}(v-\hat{u}_{B})\|_{\ell^{2}}^{2}+\|P_{N_{0}}^{\perp}(v-\hat{u}_{B})\|_{\ell^{2}}^{2}.$ (23)
We can bound the first term on the right-hand side of (23) by means of Lemma 2.9 and get, for any $B\in\mathcal{B}$, that
$\|P_{N_{0}}(v-\hat{u}_{B})\|_{\ell^{2}}^{2}$ $\leq\frac{1}{c_{\rm LB}}\|\Sigma_{\varepsilon}^{-1/2}ABP_{N_{0}}(v-\hat{u}_{B})\|_{Y}^{2}$
$\leq\frac{2}{c_{\rm LB}}\|\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\|_{Y}^{2}+\frac{2L_{F^{\prime}}}{c_{\rm LB}}\|P_{N_{0}}^{\perp}(v-\hat{u}_{B})\|_{\ell^{2}}^{2},$
where the constant $c_{\rm LB}$ depends on $N_{0}$, which is nevertheless independent of $B$. Collecting now the terms and observing that (22) trivially holds with the larger constant $C=\max\left\{2c_{\rm UB}(M),\frac{4}{c_{\rm LB}}\left(1+\frac{2L_{F^{\prime}}}{c_{\rm LB}}\right)^{-1}\right\}$, by (23) we obtain
$\|v-\hat{u}_{B}\|_{\ell^{2}}^{2}\leq\frac{2\|\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\|_{Y}^{2}}{c_{\rm LB}}+\left(1+\frac{2L_{F^{\prime}}}{c_{\rm LB}}\right)\|P_{N_{0}}^{\perp}(v-\hat{u}_{B})\|_{\ell^{2}}^{2}$
$\leq\frac{2\|\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\|_{Y}^{2}}{c_{\rm LB}}+C\left(1+\frac{2L_{F^{\prime}}}{c_{\rm LB}}\right)\left(J_{B}(v)-J_{B}(\hat{u}_{B})-\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}AB(v-\hat{u}_{B})\|_{Y}^{2}\right).$
Since $\frac{2}{c_{\rm LB}}-\frac{C}{2}\left(1+\frac{2L_{F^{\prime}}}{c_{\rm LB}}\right)\leq 0$ we have that (19) holds with
$\tilde{c}_{\rm ST}=C\left(1+\frac{2L_{F^{\prime}}}{c_{\rm LB}}\right)=\max\left\{2c_{\rm UB}(M)\left(1+\frac{2L_{F^{\prime}}}{c_{\rm LB}}\right),\frac{4}{c_{\rm LB}}\right\}.$ (24)
This completes the proof. ∎
Now the uniqueness of the minimizer $\hat{u}_{B}$ follows directly from Proposition 2.10.
Proof of Thm. 2.3 (uniqueness).
Let $\hat{u}$ and $\hat{u}^{\prime}$ be two minimizers of $J_{B}$. By Proposition 2.10, choosing $v=\hat{u}^{\prime}$ and $M=0$ (indeed, $J_{B}(\hat{u})=J_{B}(\hat{u}^{\prime})\leq J_{B}(0)=0$) we would have
$\|\hat{u}-\hat{u}^{\prime}\|_{\ell^{2}}^{2}\leq\tilde{c}_{\rm ST}(J_{B}(\hat{u})-J_{B}(\hat{u}^{\prime}))=0.$
Finally, we obtain a stability estimate for the perturbation of $\hat{u}_{B}$ with respect to modifications of $B$ within the class $\mathcal{B}$.
Proof of Thm. 2.5.
Notice that, thanks to the uniform bound on the minimizers (15) it holds that, independently of $B_{1},B_{2}\in\mathcal{B}$,
$J_{B_{2}}(\hat{u}_{1})$ $=$ $\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}AB_{2}\hat{u}_{1}\|_{Y}^{2}-\langle y,\Sigma_{\varepsilon}^{-1}AB_{2}\hat{u}_{1}\rangle_{Y}+\|\hat{u}_{1}\|_{\ell^{1}}$ (25)
$\leq$ $\frac{1}{2}L_{F^{\prime}}c_{\rm UB}+Q\|\Sigma_{\varepsilon}^{-1}A\||\mathcal{B}|c_{\rm UB}+c_{\rm UB}.$
Then, we can apply Proposition 2.10 to $J_{B_{2}}$ for the choice $v=\hat{u}_{1}$ with a $M$-level set specified by the upper bound in (25). We obtain
$\|\hat{u}_{1}-\hat{u}_{2}\|_{\ell^{2}}^{2}\leq\tilde{c}_{\rm ST}(J_{B_{2}}(\hat{u}_{1})-J_{B_{2}}(\hat{u}_{2})),$
where the explicit expression of the constant $\tilde{c}_{\rm ST}$ is given in (24). Now, since $J_{B_{1}}(\hat{u}_{1})\leq J_{B_{1}}(\hat{u}_{2})$, we have
$\|\hat{u}_{1}-\hat{u}_{2}\|_{\ell^{2}}^{2}$ $\leq\tilde{c}_{\rm ST}\big{(}J_{B_{2}}(\hat{u}_{1})-J_{B_{2}}(\hat{u}_{2})+J_{B_{1}}(\hat{u}_{2})-J_{B_{1}}(\hat{u}_{1})\big{)}$
$\leq\tilde{c}_{\rm ST}\big{(}|J_{B_{2}}(\hat{u}_{1})-J_{B_{1}}(\hat{u}_{1})|+|J_{B_{1}}(\hat{u}_{2})-J_{B_{2}}(\hat{u}_{2})|\big{)}.$
Consider now each term $|J_{B_{1}}(\hat{u}_{i})-J_{B_{2}}(\hat{u}_{i})|$ with $i=1,2$. Since $\|\hat{u}_{i}\|_{\ell^{2}}\leq c_{\rm UB}$, we have
$|J_{B_{1}}(\hat{u}_{i})-J_{B_{2}}(\hat{u}_{i})|$ $\leq\big{|}\|\Sigma_{\varepsilon}^{-1/2}AB_{1}\hat{u}_{i}\|_{Y}^{2}-\|\Sigma_{\varepsilon}^{-1/2}AB_{2}\hat{u}_{i}\|_{Y}^{2}\big{|}+\big{|}\langle y,\Sigma_{\varepsilon}^{-1}A(B_{1}-B_{2})\hat{u}_{i}\rangle_{Y}\big{|}$
$\leq\big{|}\langle\Sigma_{\varepsilon}^{-1}A(B_{1}-B_{2})\hat{u}_{i},A(B_{1}+B_{2})\hat{u}_{i}\rangle_{Y}\big{|}+\big{|}\langle\big{(}\Sigma_{\varepsilon}^{-1}A\big{)}^{*}y,(B_{1}-B_{2})\hat{u}_{i}\rangle\big{|}$
$\leq\Big{(}\|\Sigma_{\varepsilon}^{-1}A\|\|A\|\|B_{1}+B_{2}\|\|\hat{u}_{i}\|_{\ell^{2}}^{2}+\|\Sigma_{\varepsilon}^{-1}A\|\|y\|\|\hat{u}_{i}\|_{\ell^{2}}\Big{)}\|B_{1}-B_{2}\|$
$\leq(2|\mathcal{B}|\|\Sigma_{\varepsilon}^{-1}A\|\|A\|c_{\rm UB}^{2}+\|\Sigma_{\varepsilon}^{-1}A\|Qc_{\rm UB})\|B_{1}-B_{2}\|=C\|B_{1}-B_{2}\|,$
where $C=\|\Sigma_{\varepsilon}^{-1}A\|c_{\rm UB}(2|\mathcal{B}|\|A\|c_{\rm UB}+Q).$ Thus, we obtain
$\|\hat{u}_{1}-\hat{u}_{2}\|_{\ell^{2}}\leq\left(2C\tilde{c}_{\rm ST}\|B_{1}-B_{2}\|\right)^{1/2},$
and, in conclusion, since $\hat{x}_{i}=B_{i}\hat{u}_{i}$, it follows that
$\|\hat{x}_{1}-\hat{x}_{2}\|_{\ell^{2}}$ $\leq\|B_{1}-B_{2}\|\|\hat{u}_{1}\|_{\ell^{2}}+\|B_{2}\|\|\hat{u}_{1}-\hat{u}_{2}\|_{\ell^{2}}$
$\leq\big{(}\|B_{1}-B_{2}\|^{1/2}\|\hat{u_{1}}\|+\|B_{2}\|(2C\tilde{c}_{\rm ST})^{1/2})\|B_{1}-B_{2}\|^{1/2}$
$\leq\big{(}(2|\mathcal{B}|)^{1/2}c_{\rm UB}+|\mathcal{B}|(2\tilde{c}_{\rm ST}C)^{1/2}\big{)}\|B_{1}-B_{2}\|^{1/2},$
which concludes the proof with the choice $c_{\rm ST}=(2|\mathcal{B}|)^{1/2}c_{\rm UB}+|\mathcal{B}|(2\tilde{c}_{\rm ST}C)^{1/2}.$
3 Statistical learning framework
As shown in the previous section, for every choice of $B\in\mathcal{B}$ and any noisy output $y\in Y$, we can associate a solution $x=R_{B}(y)\in X$ of the inverse problem $Ax=y$, which depends on $B$. As discussed in Section 1 (see in particular Assumption 1.1), the output $y$ and hence the solution $R_{B}(y)$ are random variables, and the optimal $B^{\star}$ is defined as the minimizer of the expected risk $L$ defined by (4). Since $L$ depends on the unknown distribution of the pair $(x,y)$, $B^{\star}$ is estimated by its empirical version $\widehat{B}$, given by (7). In this section, we provide a finite sample bound on the discrepancy
$L(\widehat{B})-L(B^{\star}),$
under the assumption that both $x$ and $\varepsilon$ are bounded random variables, and $\mathcal{B}$ is compact. In fact, boundedness is assumed for convenience; for similar techniques with sub-Gaussian random variables, see e.g. [38].
Assumption 3.1.
There exists $Q_{0}>0$ such that $\left\|x\right\|_{X}\leq Q_{0}$ and $\left\|\varepsilon\right\|_{Y}\leq Q_{0}$ almost surely.
Assumption 3.1 directly implies a bound $\left\|y\right\|_{Y}\leq Q=(\left\|A\right\|+1)Q_{0}$, which was used in Section 2.
The following result shows the existence of $B^{\star}$ and $\widehat{B}$, and provides a finite sample bound on $L(\widehat{B})-L(B^{\star})$ in probability. It is proved analogously to Proposition 4 in [16] (see also [4, Lemma A.5]).
We recall that, for any $r>0$, $\mathcal{N}(\mathcal{B},r)$ denotes the covering number of $\mathcal{B}$, i.e., the minimum number of balls of radius $r$ (in the strong operator norm on $\mathcal{L}(\ell^{2},X)$) whose union contains $\mathcal{B}$.
Theorem 3.2.
Let Assumptions 1.1, 2.1, 2.2, 2.4, and 3.1 hold. There exist a minimizer $B^{\star}$ of $L$ and, with probability $1$, a minimizer $\widehat{B}$ of $\widehat{L}$ over $\mathcal{B}$. Furthermore, for all $\eta>0$,
$\mathbb{P}_{{\mathbf{z}}\sim\rho^{m}}\left[L(\widehat{B})-L(B^{\star})\leq\eta\right]\geq 1-2\mathcal{N}\left(\mathcal{B},C\eta^{2}\right)e^{-C^{\prime}m\eta^{2}}$ (26)
for some $C,C^{\prime}>0$ depending only on $A$, $\Sigma_{\varepsilon}$, $Q_{0}$ and $\mathcal{B}$.
In the above statement, the quantity $\widehat{B}$ depends on ${\mathbf{z}}$ and is therefore a random variable. For ease and convenience, in the rest of the section, we denote by $\mathbb{P}$ and $\mathbb{E}$ the probability and the expected value computed with respect to the sample ${\mathbf{z}}$, respectively.
Proof.
Consider two different elements $B_{1},B_{2}\in\mathcal{B}$. Then, $\rho$-a.e. in $X\times Y$
$\left|\|R_{B_{1}}(y)-x\|_{X}^{2}-\|R_{B_{2}}(y)-x\|_{X}^{2}\right|$ $=\left|\langle R_{B_{1}}(y)-R_{B_{2}}(y),R_{B_{1}}(y)-x+R_{B_{2}}(y)-x\rangle\right|$
$\leq\|R_{B_{1}}(y)-R_{B_{2}}(y)\|_{X}(\|R_{B_{1}}(y)\|_{X}+\|R_{B_{2}}(y)\|_{X}+2Q_{0})$
$\leq 2(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}\|B_{1}-B_{2}\|^{1/2},$
here we have used Theorem 2.5 and (15). By integrating with respect to the probability distribution $\rho$, the above bound holds for $L$. Indeed,
$|L(B_{1})-L(B_{2})|$ $=$ $\left|\mathbb{E}[\|R_{B_{1}}(y)-x\|_{X}^{2}]-\mathbb{E}[\|R_{B_{2}}(y)-x\|_{X}^{2}]\right|$ (27)
$\leq$ $\mathbb{E}\left[|\|R_{B_{1}}(y)-x\|_{X}^{2}-\|R_{B_{2}}(y)-x\|_{X}^{2}|\right]$
$\leq$ $2(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}\|B_{1}-B_{2}\|^{1/2},$
and, by replacing $\rho$ with its empirical counterpart, with probability $1$,
$\begin{split}|\widehat{L}(B_{1})-\widehat{L}(B_{2})|&=\left|\frac{1}{m}\sum_{j=1}^{m}\|R_{B_{1}}(y_{j})-x_{j}\|_{X}^{2}-\frac{1}{m}\sum_{j=1}^{m}\|R_{B_{2}}(y_{j})-x_{j}\|_{X}^{2}\right|\\ &\leq\frac{1}{m}\sum_{j=1}^{m}\left|\|R_{B_{1}}(y_{j})-x_{j}\|_{X}^{2}-\|R_{B_{2}}(y_{j})-x_{j}\|_{X}^{2}\right|\\ &\leq 2(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}\|B_{1}-B_{2}\|^{1/2}.\end{split}$ (28)
Since both $L$ and $\widehat{L}$ are continuous functionals on $\mathcal{B}$ and $\mathcal{B}$ is compact with respect to the operator topology, the corresponding minimizers $B^{\star}$ and $\widehat{B}$ over $\mathcal{B}$ exist (almost surely for $\widehat{B}$).
Next, we notice that
$\sup_{B\in\mathcal{B}}|\widehat{L}(B)-L(B)|\leq\frac{\eta}{2}\quad\Rightarrow\quad L(\widehat{B})-\widehat{L}(\widehat{B})\leq\frac{\eta}{2}\quad\text{and}\quad\widehat{L}(B^{\star})-L(B^{\star})\leq\frac{\eta}{2},$
which ultimately implies that
$0\leq L(\widehat{B})-L(B^{\star})=\left(L(\widehat{B})-\widehat{L}(\widehat{B})\right)+\left(\widehat{L}(\widehat{B})-\widehat{L}(B^{\star})\right)+\left(\widehat{L}(B^{\star})-L(B^{\star})\right)\leq\eta,$
since the central difference is negative by the definition of $\widehat{B}$. Thus,
$\mathbb{P}\left[L(\widehat{B})-L(B^{\star})\leq\eta\right]\geq\mathbb{P}\left[\sup_{B\in\mathcal{B}}|\widehat{L}(B)-L(B)|\leq\frac{\eta}{2}\right].$
We now provide a lower bound for the latter term. In view of (27) and (28), by using the reverse triangle inequality, for every $B_{1},B_{2}\in\mathcal{B}$,
$\left||\widehat{L}(B_{1})-L(B_{1})|-|\widehat{L}(B_{2})-L(B_{2})|\right|$ $\leq|\widehat{L}(B_{1})-\widehat{L}(B_{2})|+|L(B_{1})-L(B_{2})|$
$\leq 4(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}\|B_{1}-B_{2}\|^{1/2}.$
Let now $N=\mathcal{N}\left(\mathcal{B},\left(\frac{\eta}{16(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}}\right)^{2}\right)$ and consider a discrete set $B_{1},\ldots,B_{N}$ such that the balls $\mathcal{B}_{k}$ centered at $B_{k}$ with radius $r=\left(\frac{\eta}{16(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}}\right)^{2}$ cover the entire $\mathcal{B}$. In each ball $\mathcal{B}_{k}$, for every $B\in\mathcal{B}_{k}$ it holds
$\left||\widehat{L}(B)-L(B)|-|\widehat{L}(B_{k})-L(B_{k})|\right|\leq 4(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}\|B-B_{k}\|^{1/2}\leq\frac{\eta}{4}.$
Therefore, the event $|\widehat{L}(B)-L(B)|>\frac{\eta}{2}$ implies the event $|\widehat{L}(B_{k})-L(B_{k})|>\frac{\eta}{4}$, and a bound (in probability) of this term can be provided by standard concentration results. Indeed, $\widehat{L}(B_{k})$ is the sample average of $m$ realizations of the random variable $\|R_{B_{k}}(y)-x\|_{X}^{2}$, whose expectation is $L(B_{k})$. Moreover, such random variable is bounded by $(|\mathcal{B}|c_{\rm UB}+Q_{0})^{2}$ by assumption, and therefore via Hoeffding’s inequality
$\mathbb{P}\left[{\sup_{B\in\mathcal{B}_{k}}}|\widehat{L}(B)-L(B)|>\frac{\eta}{2}\right]$ $\leq\mathbb{P}\left[|\widehat{L}(B_{k})-L(B_{k})|>\frac{\eta}{4}\right]\leq 2e^{-\frac{m\eta^{2}}{8(|\mathcal{B}|c_{\rm UB}+Q_{0})^{4}}}.$
Notice that this inequality holds uniformly in $k$. Finally, we obtain
$\begin{split}\mathbb{P}\left[\sup_{B\in\mathcal{B}}|\widehat{L}(B)-L(B)|\leq\frac{\eta}{2}\right]&=1-\mathbb{P}\left[\sup_{B\in\mathcal{B}}|\widehat{L}(B)-L(B)|>\frac{\eta}{2}\right]\\ &\geq 1-\sum_{k=1}^{N}\mathbb{P}\left[\sup_{B\in\mathcal{B}_{k}}|\widehat{L}(B)-L(B)|>\eta\right]\\ &\geq 1-2Ne^{-\frac{m\eta^{2}}{8(|\mathcal{B}|c_{\rm UB}+Q_{0})^{4}}}.\end{split}$
This concludes the proof. ∎
We now provide a more explicit expression for the sample error estimate under the assumption that the covering numbers of the set $\mathcal{B}$ have a specific decay rate. It is possible to obtain similar bounds, for example, whenever the singular values of the compact embedding of $\mathcal{B}$ into its ambient space show a polynomial decay.
Corollary 3.3.
Let Assumptions 1.1, 2.1, 2.2, 2.4, and 3.1 hold. Assume further that
$\log(\mathcal{N}(\mathcal{B},r))\leq Cr^{-1/s},$ (29)
holds for some constants $C,s>0$. Suppose that $\tau>0$. Then, for sufficiently large values of $m$, there holds
$\mathbb{P}\left[L(\hat{B})-L(B^{\star})\leq\left(\frac{\alpha_{1}+\alpha_{2}\sqrt{\tau}}{\sqrt{m}}\right)^{1-\frac{1}{1+s}}\right]\geq 1-e^{-\tau}$ (30)
for some $\alpha_{1},\alpha_{2}>0$ depending only on $A$, $\mathcal{B}$, $\Sigma_{\varepsilon}$, and $Q_{0}$.
Proof.
By (26), for $\eta\in(0,1]$ and $m$ sufficiently large (namely, such that $a_{2}m\eta^{2+2/s}\geq a_{1}$) we have
$\mathbb{P}\left[L(\widehat{B})-L(B^{\star})\leq\eta\right]\geq 1-e^{-\eta^{-2/s}(a_{2}m\eta^{2+2/s}-a_{1})}\geq 1-e^{-(a_{2}m\eta^{2+2/s}-a_{1})}=1-e^{-\tau},$ (31)
being $a_{1}=C\bigl{(}16(|\mathcal{B}|c_{\rm UB}+Q_{0})c_{\rm ST}\bigr{)}^{2/s}$, $a_{2}=(8(|\mathcal{B}|c_{\rm UB}+Q_{0})^{4})^{-1}$ and $\tau=a_{2}m\eta^{2+2/s}-a_{1}.$ We can rephrase this relationship by expressing $\eta$ as a function of $\tau$ and $m$:
$1-e^{-\tau}\leq\left[L(\widehat{B})-L(B^{\star})\leq\left(\frac{a_{1}+\tau}{a_{2}m}\right)^{\frac{1}{2+2/s}}\right]\leq\mathbb{P}\left[L(\hat{B})-L(B^{\star})\leq\left(\frac{\alpha_{1}+\alpha_{2}\sqrt{\tau}}{\sqrt{m}}\right)^{1-\frac{1}{1+s}}\right],$
where $\alpha_{1}=\sqrt{a_{1}/a_{2}}$ and $\alpha_{2}=\sqrt{1/a_{2}}$, as $\sqrt{a_{1}+\tau}\leq\sqrt{a_{1}}+\sqrt{\tau}$. This concludes the proof. ∎
Note that, besides the bound in probability (30), we can also provide a bound in expectation for the excess risk. Indeed, by (31),
$\mathbb{P}\left[L(\widehat{B})-L(B^{\star})<\eta\right]\leq\min\big{\{}1,e^{a_{1}\eta^{-2/s}-a_{2}m\eta^{2}}\big{\}},$
and by the tail integral formula (notice that $e^{a_{1}\eta^{-2/s}-a_{2}m\eta^{2}}=1$ for $\eta=\hat{\eta}=k_{1}m^{-\frac{s}{2(s+1)}}$)
[…]
which allows us to conclude that, for sufficiently large $m$,
$\mathbb{E}\big{[}L(\hat{B})-L(B^{\star})\big{]}\lesssim m^{-\frac{s}{2(s+1)}}.$ (32)
The bound (32) should be compared with the one in [38, Corollary 7.2] setting $\alpha=\frac{1}{2}$ and $q=2$, namely
$\mathbb{E}\big{[}L(\hat{B})-L(B^{\star})\big{]}\lesssim m^{-\frac{1}{2}}$ (33)
provided that $s>1$, compare with the assumptions of [38, Proposition 7.3]. The rate in (32) is worse than the rate in (33) obtained via a chaining argument. On the other side, the bound in Corollary 3.3 holds in probability, whereas the bound (33) holds only in expectation.
4 On the choice of the parameter class $\mathcal{B}$
4.1 Compact perturbations of a known operator
We first consider an example of a class $\mathcal{B}$ consisting of certain compact perturbations of a reference operator $B_{0}$. Under rather general assumptions, we can provide an estimate of the covering numbers of $\mathcal{B}$. The proofs of this section are standard, and are postponed to Appendix D.
Assume that $A$ is injective. We fix an injective operator $B_{0}\colon\ell^{2}\rightarrow X$ and two compact operators $E_{1},E_{2}\colon\ell^{2}\to\ell^{2}$ such that $\|B_{0}\|_{\mathcal{L}(\ell^{2},X)}\leq 1$ and $\|E_{i}\|_{\mathcal{L}(\ell^{2})}\leq 1$ for $i=1,2$ (here, $\mathcal{L}(X):=\mathcal{L}(X,X)$). For every finite set $I\subset\mathbb{N}$, let $c_{I}>0$. We define
=$\mathcal{B}=\{B_{0}(\operatorname{Id}+K):K\in\mathcal{H}\},$ (34)
where the set of operators $\mathcal{H}$ is defined as
=$\mathcal{H}=\{K=E_{1}TE_{2}:T\in\mathcal{L}(\ell^{2}),\;\|T\|_{\mathcal{L}(\ell^{2})}\leq 1\;\text{and}\\ \|(\operatorname{Id}+K)u\|_{\ell^{2}}\geq c_{I}\|u\|_{\ell^{2}}\;\text{ for every finite $I\subset\mathbb{N}$ and $u\in\ell^{2}_{I}$}\},$ (35)
where $\ell^{2}_{I}:=\operatorname{span}\{e_{i}:i\in I\}$.
Lemma 4.1.
The set $\mathcal{B}$ defined in (34)-(35) satisfies Assumption 2.4.
The theoretical results of Section 3 rely on the covering numbers of $\mathcal{B}$. We now derive an estimate of the form (29) in the case when $\mathcal{B}$ is given by (34). For the sake of clarity, we indicate the norm used to perform the covering explicitly.
Proposition 4.2.
Let $\mathcal{B}$ be defined as in (34)-(35). Suppose that $E_{1}=E_{2}=E$, being $E\in\mathcal{L}(\ell^{2})$ a positive operator whose eigenvalues $\lambda_{n}$ satisfy
$\lambda_{n}\leq cn^{-s},\qquad n\in\mathbb{N}$
for some $c,s>0$. Take $s^{\prime}<s$. Then there exists $C>0$ such that
$\log\mathcal{N}(\mathcal{B},r;\|\cdot\|_{\mathcal{L}(\ell^{2},X)})\leq Cr^{-\frac{2s+1}{2ss^{\prime}}},\qquad r>0.$ (36)
With this estimate on the covering numbers, it is possible to apply Corollary 3.3 and obtain a finite sample bound in the case when $\mathcal{B}$ is defined as in (34).
4.2 Learning the mother wavelet
We consider here the problem of learning the mother wavelet. In other words, the set $\mathcal{B}$ will consist of synthesis operators associated to a family of wavelets.
Consider the Hilbert space $X=L^{2}(\mathbb{R}^{d})$ and, for $\psi\in L^{2}(\mathbb{R}^{d})$, define
$\psi_{j,k}(x)=2^{\frac{jd}{2}}\psi(2^{j}x-k),\qquad\text{for a.e.\ $x\in\mathbb{R}^{d}$, $j\in\mathbb{Z}$, $k\in\mathbb{Z}^{d}$.}$
For $f\in L^{2}(\mathbb{R}^{d})\cap L^{1}(\mathbb{R}^{d})$ (the space $L^{1}(\mathbb{R}^{d})$ is added just to simplify the exposition, thanks to the continuity of $\hat{f}$), define
[…]
and consider
$W=\{f\in L^{2}(\mathbb{R}^{d})\cap L^{1}(\mathbb{R}^{d}):\|f\|_{W}<+\infty\}.$
It is easy to verify that $\|\cdot\|_{W}$ is a norm on $W$. It is well known that, for $\psi\in W$, the family $\{\psi_{j,k}\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{d}}$ is a Bessel sequence of $L^{2}(\mathbb{R}^{d})$, namely, the synthesis operator
$B_{\psi}\colon\ell^{2}(\mathbb{Z}\times\mathbb{Z}^{d})\to L^{2}(\mathbb{R}^{d}),\qquad(c_{j,k})\mapsto\sum_{j,k}c_{j,k}\psi_{j,k},$
is well defined and bounded (see, e.g., [17, Section 3.3] and [28, Theorem 3]).
We now construct a compact family of “mother wavelets” $\psi$ in $W$. (Note that the term mother wavelet here is used even though the family $\psi_{j,k}$ need not be a frame of $L^{2}(\mathbb{R}^{d})$.)
Lemma 4.3.
Let $\Psi\subseteq W$ be a compact set (with respect to $\|\cdot\|_{W}$) and $a>0$. Set
$\Psi_{a}=\{\psi\in\Psi:\|B_{\psi}x\|_{L^{2}(\mathbb{R}^{d})}\geq a\|x\|_{\ell^{2}}\ \forall x\in\ell^{2}\}.$
Then
$\mathcal{B}_{\Psi_{a}}=\{B_{\psi}:\psi\in\Psi_{a}\}$
is a compact subset of $\mathcal{L}(\ell^{2},L^{2}(\mathbb{R}^{d}))$ with respect to the operator norm, and
$\mathcal{N}\left(\mathcal{B}_{\Psi_{a}},\delta;\|\cdot\|_{\mathcal{L}(\ell^{2},L^{2}(\mathbb{R}^{d}))}\right)\leq\mathcal{N}\left(\Psi,(2\pi)^{\frac{3}{2}d}\,\delta;\|\cdot\|_{W}\right),\qquad\delta>0.$
Proof.
By the estimates in the proof of Theorem 3 in [28], we have
$\sum_{j,k}|\langle f,\psi_{j,k}\rangle|^{2}\leq(2\pi)^{-3d}\|\psi\|_{W}^{2}\|f\|^{2}_{L^{2}(\mathbb{R}^{d})},\qquad\psi\in W,\,f\in L^{2}(\mathbb{R}^{d}).$
In other words, we have the following bound on the norm of the analysis operator: $\|B^{*}_{\psi}\|\leq(2\pi)^{-\frac{3}{2}d}\|\psi\|_{W}.$ As a consequence, we have
$\|B_{\psi}\|\leq(2\pi)^{-\frac{3}{2}d}\|\psi\|_{W},\qquad\psi\in W.$
In other words, the linear map
$\zeta\colon W\to\mathcal{L}(\ell^{2},L^{2}(\mathbb{R}^{d})),\qquad\psi\mapsto B_{\psi}$
is linear and bounded, with $\|\zeta\|\leq(2\pi)^{-\frac{3}{2}d}$.
Since $\zeta$ is bounded and $\Psi$ is compact, we have that $\{B_{\psi}:\psi\in\Psi\}=\zeta(\Psi)$ is compact. Thus, $\mathcal{B}_{\Psi}$ is the intersection of a compact set and a closed set, and so it is compact.
Finally, the estimate on the covering numbers of $\mathcal{B}_{\Psi_{a}}$ immediately follows:
$\mathcal{N}\left(\mathcal{B}_{\Psi_{a}},\delta\right)\leq\mathcal{N}\left(\zeta(\Psi),\delta\right)\leq\mathcal{N}\left(\Psi,\|\zeta\|^{-1}\delta\right)\leq\mathcal{N}\left(\Psi,(2\pi)^{\frac{3}{2}d}\,\delta\right).$
We can now combine this result with Corollary 3.3 and obtain the following corollary, in which we compare the loss corresponding to the optimal mother wavelet $\psi^{*}$ with the loss corresponding to the mother wavelet $\widehat{\psi}$ obtained by minimizing the empirical risk.
Corollary 4.4.
Consider the settings of Corollary 3.3 and of Lemma 4.3. Suppose that
$\log\mathcal{N}(\Psi,r;\|\cdot\|_{W})\leq Cr^{-1/s},\qquad r>0,$
for some $C,s>0$. There exist $\alpha_{1},\alpha_{2}>0$ such that the following is true. Take $\tau>0$ and
$\psi^{*}\in\operatorname*{arg\,min}_{\psi\in\Psi_{a}}L(B_{\psi}),\qquad\widehat{\psi}\in\operatorname*{arg\,min}_{\psi\in\Psi_{a}}\widehat{L}(B_{\psi}).$
Then, with probability larger than $1-e^{-\tau}$, we have
$L(B_{\widehat{\psi}})-L(B_{\psi^{*}})\leq\left(\frac{\alpha_{1}+\alpha_{2}\sqrt{\tau}}{\sqrt{m}}\right)^{1-\frac{1}{1+s}}.$
Proof.
By Lemma 4.3, we have
$\log\mathcal{N}\left(\mathcal{B}_{\Psi_{a}},r\right)\leq\log\mathcal{N}\left(\Psi,(2\pi)^{\frac{3}{2}d}\,r\right)\leq C(2\pi)^{-\frac{3d}{2s}}r^{-1/s}.$
The result is now a direct consequence of Corollary 3.3. ∎
Appendix A Finite-dimensional setting
In order to derive a more interpretable expression for the mininization problem (2), let us first suppose that, on top of Assumption 1.1, the covariance operator $\Sigma_{\varepsilon}\colon Y\rightarrow Y$ is also surjective. Notice in particular that this can only hold if the space $Y$ is finite-dimensional, for otherwise it would be impossible for $\Sigma_{\varepsilon}$ to be both compact and invertible. In this case, we can add the term $\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}y\|_{Y}^{2}$ in (2) (which is irrelevant for optimization purposes) and get
$\hat{x}_{B}=B\hat{u}_{B},\quad\hat{u}_{B}=\operatorname*{arg\,min}_{u\in\ell^{1}}\Big{\{}\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}(ABu-y)\|_{Y}^{2}+\|u\|_{\ell^{1}}\Big{\}}.$ (37)
When $\Sigma_{\varepsilon}=\sigma^{2}I$, this reduces to the familiar form of Lasso regression with $\ell^{1}$ regularization. In this context, it is easier to interpret (37) as a regularization of the inverse problem (1), which promotes the sparsity of the solution with respect to the frame associated with the operator $B$. In particular, the expression in (37) is known as the synthesis formulation of such problem, since the synthesis operator $B\colon\ell^{2}\rightarrow X$ appears in it, and the minimization takes place in the whole space of coefficients $\ell^{2}$. An alternative problem, which is in general not equivalent to (37) is the analysis formulation of the sparsity-promoting regularization: for a bounded linear operator $C\colon X\rightarrow\ell^{2}$, let
$\tilde{x}_{C}=\operatorname*{arg\,min}_{x\in X}\Big{\{}\frac{1}{2}\|\Sigma_{\varepsilon}^{-1/2}(Ax-y)\|_{Y}^{2}+\|Cx\|_{\ell^{1}}\Big{\}}.$ (38)
The equivalence of (37) and (38) holds if $B$ is invertible, and $C=B^{-1}$: in that case, it holds that $\tilde{x}_{B^{-1}}=\hat{x}_{B}$. The task of learning an optimal sparsity-promoting regularizer might have been expressed in terms of the analysis operator $C$ but, for theoretical purposes, we preferred considering the synthesis formulation.
Appendix B White noise
As we mentioned in Section 1, the assumption on $\Sigma_{\varepsilon}$ does not allow us, in principle, to consider white noise in infinite dimensions, for which the covariance $\Sigma_{\varepsilon}=\mathrm{Id}$ is not trace-class, but only bounded. However, this situation can be considered by embedding $Y$ into a larger space. Instead of providing a general and abstract construction, we prefer to discuss two particular examples. For more details, see [4, Section A.1].
Example B.1.
Let $\Omega$ be a connected bounded open subset of $\mathbb{R}^{d}$, and $\{\varepsilon_{y}\}_{y\in L^{2}(\Omega)}$ be a Hilbert random process such that
$\mathbb{E}{[\varepsilon_{y}]}=0\qquad\mathbb{E}{[\varepsilon_{y}\varepsilon_{y^{\prime}}]}=\langle y,y^{\prime}\rangle_{L^{2}(\Omega)},\qquad y,y^{\prime}\in L^{2}(\Omega),$
which is the classical model for white noise, since the covariance operator of $\varepsilon$ is the identity. Assume that $\Omega$ satisfies a cone condition and fix $s>d/2$. Then the Rellich-Kondrachov theorem [1, Theorem 6.3] implies that the Sobolev space $H^{s}(\Omega)$ is embedded into $C_{b}(\Omega)$, the space of bounded continuous functions on $\Omega$, and the canonical inclusion $\iota\colon H^{s}(\Omega)\to L^{2}(\Omega)$ is a compact operator. The above embedding implies that $H^{s}(\Omega)$ is a reproducing kernel Hilbert space with bounded reproducing kernel $K\colon\Omega\times\Omega\to\mathbb{R}$. An easy calculation shows that
[…]
This fact implies that in the Gelfand triple
[…]
both $\iota$ and $\iota^{*}$ are Hilbert-Schmidt operators. Hence, it is possible to define a random variable $\widehat{\varepsilon}$ taking values in $H^{-s}(\Omega)$ such that
$\langle\widehat{\varepsilon},y\rangle=\varepsilon_{\iota(y)},\qquad y\in H^{s}(\Omega).$
It is immediate to check that
$\mathbb{E}{[\langle\widehat{\varepsilon},y\rangle_{H^{-s},H^{s}}\ \langle\widehat{\varepsilon},y^{\prime}\rangle_{H^{-s},H^{s}}]}=\langle\iota(y),\iota(y^{\prime})\rangle_{L^{2}(\Omega)},$
so that the covariance operator of $\widehat{\varepsilon}$ is $\iota^{*}\circ\iota$, which is a trace class operator. Thus, $\widehat{\varepsilon}$ is a square-integrable random variable in $H^{-s}(\Omega)$. To apply the results of our paper, it is enough to set $Y=H^{-s}(\Omega)$ and to lift the inverse problem (1) to $H^{-s}(\Omega)$:
$\iota^{*}(y)=(\iota^{*}\circ A)x+\widehat{\varepsilon}.$
This requires identifying $H^{s}(\Omega)$ and $H^{-s}(\Omega)$ using the Riesz lemma. Note that this identification is not standard, since it is not compatible with the double embedding of (39). However, the intermediate space $L^{2}(\Omega)$ does not matter once $\widehat{\varepsilon}$ is defined.
In the next example, we consider the sequence space $\ell^{2}$, as a prototypical Hilbert space (once an orthonormal basis has been fixed).
Example B.2.
Let $X$ be a Hilbert space and set $Y=\ell^{2}$. Let $A\colon X\to\ell^{2}$ be a bounded and linear map, and consider the inverse problem (1)
$y=Ax+\varepsilon.$ (40)
Let $\{e_{n}\}_{n\in\mathbb{N}}$ be the canonical basis of $\ell^{2}$, defined by $(e_{n})_{i}=\delta_{i,n}$. We consider the case when $\varepsilon$ is a white Gaussian noise with variance $\sigma^{2}$, namely,
$\varepsilon=\sum_{n\in\mathbb{N}}\varepsilon_{n}e_{n},$
where the random variables $\varepsilon_{n}$ are i.i.d. scalar Gaussians with mean zero and variance $\sigma^{2}$, i.e.  $\varepsilon_{n}\sim\mathcal{N}(0,\sigma^{2})$. The above expression for $\varepsilon$ is only formal, because the series is divergent in $\ell^{2}$ with probability 1. As a consequence, (40) is not well-defined in $\ell^{2}$. However, writing (40) in components with respect to the orthonormal basis $\{e_{n}\}_{n\in\mathbb{N}}$ yields
$y_{n}=(Ax)_{n}+\varepsilon_{n},\qquad n\in\mathbb{N},$ (41)
where we wrote $Ax=\sum_{n}(Ax)_{n}e_{n}$. This is a family of well-defined scalar equations.
Let us now see how it is possible to reformulate this as a problem in a Hilbert space. Equivalently, we can rewrite (41) as
$\frac{y_{n}}{n^{s}}=\frac{(Ax)_{n}}{n^{s}}+\frac{\varepsilon_{n}}{n^{s}},\qquad n\in\mathbb{N},$ (42)
for some $s>\frac{1}{2}$. Let us introduce the embedding $\iota^{*}\colon\ell^{2}\to\ell^{2}$ defined by $\iota^{*}(e_{n})=e_{n}/n^{s}$ and the random variable
$\widehat{\varepsilon}=\sum_{n\in\mathbb{N}}\frac{\varepsilon_{n}}{n^{s}}e_{n}.$
Note that $\mathbb{E}{[\left\|\widehat{\varepsilon}\right\|^{2}_{2}]}<+\infty$, and $\widehat{\varepsilon}\in\ell^{2}$ with probability 1. Thus, we can rewrite (42) as
$\iota^{*}(y)=\iota^{*}(Ax)+\widehat{\varepsilon}.$
This equation is meaningful in $\ell^{2}$, and has the same form as the original inverse problem (40).
Appendix C Connections with Dictionary Learning and unsupervised strategies
Although our approach shares the same aim of dictionary learning, i.e., promoting the sparse representation of some ground truths by selecting a suitable synthesis operator, it is possible to outline some substantial differences. The key observation is that the optimal operator sought in dictionary learning is independent of the forward operator and the noise distribution, and depends only on the distribution of $x$. On the contrary, the optimal $\widehat{B}$ in (5) depends on both $\varepsilon$ and $A$, yielding, in general, a smaller MSE and, consequently, better statistical guarantees for the solution of the inverse problem.
Let us briefly introduce the standard dictionary learning framework [40]. If, instead of the training dataset ${\mathbf{z}}=\{(x_{j},y_{j})\}_{j=1}^{m}$ employed in (7) to discretize (5), only a collection of ground truths $\{x_{j}\}_{j=1}^{m}$ is available, i.i.d. sampled from the (unknown) marginal $\rho_{x}$, we may consider the following unsupervised technique: let $\widetilde{R}_{B}$ a sparsity-promoting regularizer of the form (2)-(3) associated with $A=\operatorname{Id}$ and without assuming the knowledge of the covariance $\Sigma_{\varepsilon}$, namely:
$\widetilde{R}_{B}(x)=B\widetilde{u}_{B}(x),\quad\widetilde{u}_{B}(x)=\operatorname*{arg\,min}_{u\in\ell^{1}}\left\{\frac{1}{2}\|Bu-x\|^{2}+\|u\|_{\ell^{1}}\right\}$
$\widehat{B}_{DL}\in\operatorname*{arg\,min}_{B\in\mathcal{B}}\left\{\frac{1}{m}\sum_{j=1}^{m}\|x_{j}-\widetilde{R}_{B}(x_{j})\|^{2}\right\}.$
This problem yields a bilevel formulation of the well-known Dictionary Learning problem [37, 42]. Our supervised strategy resembles dictionary learning, indeed, let us recall that, according to (7),
$\widehat{B}$ $\in\operatorname*{arg\,min}_{B\in\mathcal{B}}\left\{\frac{1}{m}\sum_{j=1}^{m}\|x_{j}-R_{B}(Ax_{j}+\varepsilon_{j})\|^{2}\right\},$
being $R_{B}$ as in (3). However, since $R_{B}$ is a nonlinear map, differently from our previous work on quadratic regularizers [4], it is easy to show that the two problems are not equivalent in general. In particular, while $\widehat{B}_{DL}$ is independent of the forward operator $A$ and on the noise $\varepsilon$ by construction, $\widehat{B}$ will in general depend on both. We illustrate this with a simple 1D example.
Let $\sigma^{2}=\mathbb{E}(\varepsilon^{2})>0$ denote the variance of the noise (with zero mean) and consider the 1D problem
$Ax=ax\quad\text{with }a\in(0,\sigma]$
and the regularization $B^{-1}x=bx$, $b>0$. Given an unknown $x^{\dagger}$ and data $y=Ax^{\dagger}+\varepsilon=ax^{\dagger}+\varepsilon$, the Lasso reconstruction is given by
$\hat{x}=\operatorname*{arg\,min}_{x}\frac{1}{2}|\sigma^{-1}(Ax-y)|^{2}+|bx|=\operatorname*{arg\,min}_{x}\frac{1}{2}|x-y/a|^{2}+\frac{\sigma^{2}b}{a^{2}}|x|.$
Setting $\gamma=\frac{a}{\sigma}$, this may be rewritten by using the soft-thresholding operator $S_{\lambda}$ as
$\hat{x}=S_{\frac{b}{\gamma^{2}}}\left(x^{\dagger}+\gamma^{-1}\tilde{\varepsilon}\right),$
where $\tilde{\varepsilon}=\varepsilon/\sigma$ satisfies $\mathbb{E}(\tilde{\varepsilon}^{2})=1$. Now consider the mean squared error
$\text{MSE}=\mathbb{E}_{x,\tilde{\varepsilon}}\left[\left|S_{\frac{b}{\gamma^{2}}}\left(x+\gamma^{-1}\tilde{\varepsilon}\right)-x\right|^{2}\right].$
Note that we consider the minimization of the expected risk, and not of the empirical risk, because it is more significant. For simplicity, we choose the following zero-mean independent distributions
$\mathbb{P}(x=\pm 1)=\frac{1}{2},\quad\mathbb{P}(\tilde{\varepsilon}=\pm 1)=\frac{1}{2}.$
After a series of elementary computations, we can show that
[…]
Therefore, the optimal value for the MSE is achieved at $b=\gamma=\frac{a}{\sigma}$ and depends on both $a$ and $\sigma$, namely, on both the forward operator and the noise level. Any unsupervised choice for $b$ would be independent of $a$ and $\sigma$ and give rise, in general, to larger values of the MSE.
Appendix D Proofs of Section 4.1
Proof of Lemma 4.1.
Every element $B\in\mathcal{B}$ satisfies Assumption 2.2 because $A$ is injective and $B|_{\ell^{2}_{I}}$ is injective for every finite subset $I\subset\mathbb{N}$ by (35).
It remains to show that $\mathcal{B}$ is compact. Since the map
$\mathcal{L}(\ell^{2})\ni U\longmapsto B_{0}(\operatorname{Id}+U)\in\mathcal{L}(\ell^{2},X)$ (43)
is continuous, it is enough to show that $\mathcal{H}\subset\mathcal{L}(\ell^{2})$ is compact. Write $\mathcal{H}=\mathcal{K}\cap\mathcal{C}$, where
$\mathcal{K}=\{E_{1}TE_{2}:T\in\mathcal{L}(\ell^{2}),\;\|T\|_{\mathcal{L}(\ell^{2})}\leq 1\},$
$\mathcal{C}=\{K\in\mathcal{L}(\ell^{2}):\|(\operatorname{Id}+K)u\|\geq c_{I}\|u\|\;\text{ for every finite $I\subset\mathbb{N}$ and $u\in\ell^{2}_{I}$}\}.$
The set $\mathcal{K}$ is compact by [41, Theorem 3], because $\left\{T\in\mathcal{L}(\ell^{2}):\|T\|_{\mathcal{L}(\ell^{2})}\leq 1\right\}$ is closed. The set $\mathcal{C}$ is closed, because convergence in operator norm implies pointwise convergence. Hence, $\mathcal{H}=\mathcal{K}\cap\mathcal{C}$ is compact. ∎
Proof of Proposition 4.2.
Note that
$\mathcal{N}(\mathcal{B},r;\|\cdot\|_{\mathcal{L}(\ell^{2},X)})\leq\mathcal{N}(\mathcal{H},r;\|\cdot\|_{\mathcal{L}(\ell^{2})}),$ (44)
since the map in (43) is Lipschitz with constant $1$ (recall that $\|B_{0}\|_{\mathcal{L}(\ell^{2},X)}\leq 1$). Using the notation of the proof of Lemma 4.1, we have $\mathcal{H}=\mathcal{K}\cap\mathcal{C}$, so that $\mathcal{H}\subseteq\mathcal{K}$. Thus
$\mathcal{N}(\mathcal{H},r;\|\cdot\|_{\mathcal{L}(\ell^{2})})\leq\mathcal{N}(\mathcal{K},r;\|\cdot\|_{\mathcal{L}(\ell^{2})}).$ (45)
To further bound the right-hand side of (45), we consider the singular value decomposition of $E$. Since $E$ is self-adjoint and compact, we can write its spectral decomposition in terms of its eigenvalues $\{\lambda_{n}\}$ and eigenvectors $\{e_{n}\}$, and define its rank-$N$ approximation $E_{N}$ as follows:
$E=\sum_{n=1}^{\infty}\lambda_{n}e_{n}\otimes e_{n},\quad\qquad E_{N}=\sum_{n=1}^{N}\lambda_{n}e_{n}\otimes e_{n}=P_{N}E=EP_{N},$ (46)
[…]
$\mathcal{K}_{N}=\left\{K=E_{N}TE_{N}:T\in\mathcal{L}(\ell^{2}),\ \|T\|_{\mathcal{L}(\ell^{2})}\leq 1\right\},$
and prove that the following bound holds:
$\mathcal{N}(\mathcal{K},\rho+2\lambda_{N+1};\|\cdot\|_{\mathcal{L}(\ell^{2})})\leq\mathcal{N}(\mathcal{K}_{N},\rho;\|\cdot\|_{\mathcal{L}(\ell^{2})})=:{\hat{\mathcal{N}}}.$ (47)
In order to prove (47), consider a $\rho$-covering $\{K_{1},\ldots,K_{\hat{\mathcal{N}}}\}$ of $\mathcal{K}_{N}$. For any $K=ETE\in\mathcal{K}$, let $K^{N}=E_{N}TE_{N}\in\mathcal{K}_{N}$, and let $K_{i}$ be such that $\|K^{N}-K_{i}\|\leq\rho$. Then,
$\|K-K_{i}\|_{\mathcal{L}}$ $\leq\|K-K^{N}\|_{\mathcal{L}}+\|K^{N}-K_{i}\|_{\mathcal{L}}$
$\leq\|ETE-ETE_{N}\|_{\mathcal{L}}+\|ETE_{N}-E_{N}TE_{N}\|_{\mathcal{L}}+\rho$
$\leq\|E\|_{\mathcal{L}}\|T\|_{\mathcal{L}}\|E-E_{N}\|_{\mathcal{L}}+\|E-E_{N}\|_{\mathcal{L}}\|T\|_{\mathcal{L}}\|E_{N}\|_{\mathcal{L}}+\rho$
$\leq 2\lambda_{N+1}+\rho,$
which shows that $\{K_{1},\ldots,K_{\hat{\mathcal{N}}}\}$ is a $(2\lambda_{N+1}+\rho)$-covering of $\mathcal{K}$.
In order to bound the covering numbers of $\mathcal{K}_{N}$, we claim that $\mathcal{K}_{N}\subset\mathcal{F}_{N}$, where
$\mathcal{F}_{N}=\left\{K=ETE:T\in\operatorname{HS}(\ell^{2}),\ \|T\|_{\operatorname{HS}(\ell^{2})}\leq\sqrt{N}\right\},$
and $\operatorname{HS}(\ell^{2})$ denotes the class of Hilbert-Schmidt operators from $\ell^{2}$ to $\ell^{2}$, namely, the ones for which the singular values are square-summable. In order to show this, take $K\in\mathcal{K}_{N}$, with $K=E_{N}TE_{N}$ for some $T\in\mathcal{L}(\ell^{2})$ such that $\|T\|_{\mathcal{L}(\ell^{2})}\leq 1$. Setting $T_{N}=P_{N}TP_{N}$, by (46) we have
$K=E_{N}TE_{N}=EP_{N}TP_{N}E=ET_{N}E.$
Note that $T_{N}$ is a rank-$N$ operator, thus belonging to the Hilbert-Schmidt class. Furthermore,
$\|T_{N}\|_{\operatorname{HS}(\ell^{2})}=\|P_{N}TP_{N}\|_{\operatorname{HS}(\ell^{2})}\leq\|P_{N}\|_{\operatorname{HS}(\ell^{2})}\|TP_{N}\|_{\mathcal{L}(\ell^{2})}\leq\sqrt{N},$
because $\|P_{N}\|_{\operatorname{HS}(\ell^{2})}=\sqrt{N}$, $\|TP_{N}\|_{\mathcal{L}(\ell^{2})}\leq 1$ and
$\|P_{N}TP_{N}\|^{2}_{\operatorname{HS}(\ell^{2})}=\sum_{n=1}^{\infty}\|P_{N}TP_{N}e_{n}\|^{2}_{\ell^{2}}\leq\sum_{n=1}^{\infty}\|P_{N}T\|^{2}_{\mathcal{L}(\ell^{2})}\|P_{N}e_{n}\|_{\ell^{2}}^{2}=\|P_{N}T\|^{2}_{\mathcal{L}(\ell^{2})}\|P_{N}\|_{\operatorname{HS}(\ell^{2})}^{2}.$
This shows that $K\in\mathcal{F}_{N}$, as claimed.
By a simple scaling argument, we have
$\mathcal{N}(\mathcal{K}_{N},\rho;\|\cdot\|_{\mathcal{L}})\leq\mathcal{N}(\mathcal{F}_{N},\rho;\|\cdot\|_{\mathcal{L}})=\mathcal{N}(\mathcal{F}_{1},\rho N^{-1/2};\|\cdot\|_{\mathcal{L}}).$ (48)
We finally have to estimate the covering numbers $\mathcal{N}(\mathcal{F}_{1},\varrho;\|\cdot\|_{\mathcal{L}})$. Let us define
$F_{1}=\left\{T\in\operatorname{HS}(\ell^{2}):\ \|T\|_{\operatorname{HS}(\ell^{2})}\leq 1\right\},$
which entails that $\mathcal{F}_{1}=j(F_{1})$, where $j$ is the (compact) embedding $j\colon\operatorname{HS}(\ell^{2})\rightarrow\operatorname{HS}(\ell^{2})$ defined as $j(T)=ETE$. Since $\|\cdot\|_{\mathcal{L}(\ell^{2})}\leq\|\cdot\|_{\operatorname{HS}(\ell^{2})}$, a $\varrho-$covering of $\mathcal{F}_{1}$ with respect to $\|\cdot\|_{\operatorname{HS}(\ell^{2})}$ is also a $\varrho$-covering with respect to $\|\cdot\|_{\mathcal{L}(\ell^{2})}$. Thus,
$\mathcal{N}(\mathcal{F}_{1},\varrho;\|\cdot\|_{\mathcal{L}(\ell^{2})})\leq\mathcal{N}(\mathcal{F}_{1},\varrho;\|\cdot\|_{\operatorname{HS}(\ell^{2})}).$ (49)
The quantity $\mathcal{N}(\mathcal{F}_{1},\varrho;\|\cdot\|_{\operatorname{HS}(\ell^{2})})$ is the covering number of the image of the unit sphere $F_{1}$ of the Hilbert space $\operatorname{HS}(\ell^{2})$ through the embedding $j$, and is linked to the entropy numbers $\varepsilon_{k}(j)$: indeed, according to the definitions in [14, Chapter 1], one clearly sees that
$\mathcal{N}(\mathcal{F}_{1},\varrho;\|\cdot\|_{\operatorname{HS}(\ell^{2})})\leq k\quad\iff\quad\varepsilon_{k}(j)\leq\varrho.$
In order to estimate the entropy numbers of $j$, we can rely on its singular values, thanks to [14, Theorem 3.4.2]. In view of the decay $\lambda_{n}\lesssim n^{-s}$ of the eigenvalues of $E$, following the proof of [4, Lemma A.9] we can show that the singular values of $j$ decay as $n^{-s^{\prime}}$, for any $s^{\prime}<s$. Then, by the same argument used in the proof of [4, Lemma A.8], we have that $\varepsilon_{k}(j)\lesssim(\log k)^{-s^{\prime}}$, which implies $\mathcal{N}(\mathcal{F}_{1},(\log k)^{-s^{\prime}};\|\cdot\|_{\operatorname{HS}(\ell^{2})})\leq k$ and ultimately
$\log\mathcal{N}(\mathcal{F}_{1},\varrho;\|\cdot\|_{\operatorname{HS}(\ell^{2})})\lesssim\varrho^{-1/s^{\prime}}.$ (50)
We can now conclude the proof: by (44), (45) and (47), setting $\rho=\frac{r}{2}$ and choosing $N$ such that $\lambda_{N+1}=\frac{r}{4}$ (which implies $(N+1)^{-s}\gtrsim\frac{r}{4}$, hence $N\lesssim r^{-1/s}$), we get
$\mathcal{N}(\mathcal{B},r;\|\cdot\|_{\mathcal{L}(\ell^{2},X)})\lesssim\mathcal{N}(\mathcal{K}_{N},r;\|\cdot\|_{\mathcal{L}(\ell^{2})}).$
Next, by (48), we obtain
$\mathcal{N}(\mathcal{B},r;\|\cdot\|_{\mathcal{L}(\ell^{2},X)})\lesssim\mathcal{N}(\mathcal{F}_{1},rN^{-1/2};\|\cdot\|_{\mathcal{L}(\ell^{2})}),$
and by (49) and (50) we finally obtain
$\log\mathcal{N}(\mathcal{B},r;\|\cdot\|_{\mathcal{L}(\ell^{2},X)})\lesssim\bigl{(}rN^{-1/2}\bigr{)}^{-1/s^{\prime}}\lesssim r^{-\frac{2s+1}{2ss^{\prime}}}.$
This shows (36), and the proof is concluded. ∎
Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award numbers FA8655-23-1-7083. Co-funded by the European Union (ERC, SAMPDE, 101041040, ERC, SLING, 819789, ERC, AdG project 101097198 and Next Generation EU). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. The research of LR has been funded by PNRR - M4C2 - Investimento 1.3. Partenariato Esteso PE00000013 - “FAIR - Future Artificial Intelligence Research” - Spoke 8 “Pervasive AI”, which is funded by the European Commission under the NextGeneration EU programme. The research was supported in part by the MUR Excellence Department Project awarded to Dipartimento di Matematica, Università di Genova, CUP D33C23001110001. The research by GSA, EDV and MS has been supported by the MUR grants PRIN 202244A7YL, 2022B32J5C, and P2022XT498, and by FAIR Project HAOISL 33C24000410007 funded by the European Commission under the NextGeneration EU programme. GSA, EDV, LR and MS are members of the “Istituto Nazionale di Alta Matematica”. TH and ML were supported by the Research Council of Finland (decision numbers 348504, 353094, 359182 and 359183).